OpenStudy (anonymous):

Simplify X-3[2X-4(X-1)+2]

OpenStudy (triciaal):

get rid of inside parenthesis first then outside parenthesis group like terms and simplify if possible to distribute multiply each term outside the parenthesis by each term inside the parenthesis.

OpenStudy (triciaal):

start with -4*x + -4* -1 =

OpenStudy (triciaal):

what do you get?

OpenStudy (triciaal):

-4 x + 4 next -3(2 x - 4 x + 4 + 2 ) = -3(-2 x + 6 ) = 6 x -18 x - 6 x - 18 = - 7 x - 18

OpenStudy (triciaal):